OpenStudy (anonymous):

theoretical yield of 1.42g H2 is allowed to react with 10.2g N2, producing 2.30g NH3.

OpenStudy (anonymous):

Someone help please.

OpenStudy (aaronq):

for problems involving stoichiometry, you have to convert whatever information they give you to moles (signified by the symbol "n") Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles. ------------------------------------------------- BUT first find the limiting reactant by dividing the moles of each by it's stoichiometric coefficient. Compare them now and find which has the least moles. Use the moles of limiting reactant for the calculation.